# D-Wave’s Nonlinear-Program Hybrid Solver: Description and Performance Analysis

### v2

@@ I Introduction
The emergence of quantum technology is expected to have a significant impact on a number of industries. The field of Quantum Computing (QC), which leverages the principles of quantum mechanics to process information, is constantly making advances to connect quantum processing with practical use cases. Thus, QC has advanced significantly in recent years, mainly due to the rapid development of technology and advancements in the democratization of its access [1]. As a result, QC has facilitated the development of various proofs of concept across multiple sectors, including finance [2], energy [3], and logistics [4].
Currently, there are two types of real quantum devices that we can differentiate between: gate-based quantum computers and quantum annealers. A gate-based system, on the one hand, employs basic quantum circuit operations on qubits, which are akin to the classical operations on regular bits and may be combined in any order to create algorithms. Initially, these quantum computers were created to replicate quantum physics. Later, the idea was expanded to encompass the entire processing capacity of a Turing machine. This version is often referred to as a universal quantum computer. On the other hand, a quantum annealer applies a problem-specific magnetic field after bringing a group of qubits into equal superposition. Under this magnetic field, the qubits interact and move toward the lowest-energy state, which represents the solution to an optimization problem. As a result, quantum annealers, which are specialized rather than universal systems, utilize quantum-mechanical probabilities to solve optimization problems [5].
@@
However, despite all the progress made in the field, quantum computers are still in their infancy in contrast to classical computers, which have been developed for decades and are therefore highly advanced. Thus, quantum devices are currently unable to efficiently solve real-world problems mainly because of the small number of qubits and their unstable nature, which results in errors, noise, and information loss. At this time, the number of qubits for both universal quantum computers and quantum annealers is around three orders of magnitude (thousands of qubits) [6, 7]. However, this number needs to be much greater in order for the technology to become truly useful for real industrial use cases. In any case, there are factors other than the qubit count that influence the practical capabilities of quantum computers, such as the topology layout of the qubits or the operation parallelism.
@@
In any case, it is crucial to note that the design and implementation of quantum algorithms could necessitate a high degree of subject matter expertise. This intricacy could be a hindrance for researchers who lack a sufficient background in disciplines like physics or quantum mechanics. In order to remove that barrier and make QC more accessible, various frameworks and programming languages are proposed, such as Silq [11], Eclipse Qrisp [12], or Qiskit [13]. The availability of such frameworks and languages helps to foster the building of a multidisciplinary community focused on QC and helps the field to progress toward new horizons [14].
As part of the strategy to bring QC to a wider audience, the Canadian company D-Wave Systems, a leading supplier of quantum annealers, published D-Wave-Hybrid-Framework in 2018, which, in the words of its creators, is “a general, minimal Python framework to build hybrid asynchronous decomposition samplers for QUBO problems”111https://github.com/dwavesystems/dwave-hybrid. This framework is appropriate for “developing hybrid approaches to combining quantum and classical compute resources”.
@@
As we have summarized systematically in Table I, much research has been carried out in recent years around BQM-Hybrid, DQM-Hybrid, and CQM-Hybrid. Despite this abundant scientific activity, no work has yet been published on NL-Hybrid.
Motivated by the lack of existing research, this paper focuses on describing the newly introduced NL-Hybrid solver. Furthermore, we will conduct an experiment to analyze the performance of this new solver in comparison with BQM-Hybrid, CQM-Hybrid, and D-Wave’s QPU Advantage_system6.4. For these tests, we used a benchmark composed of 45 instances, equally distributed across three combinatorial optimization problems: the Traveling Salesman Problem (TSP), the Knapsack Problem (KP), and the Maximum Cut Problem (MCP). Finally, to facilitate the use of this still unexplored solver, we will share details of the implementation to solve these three optimization problems.
@@ II-A Overview of NL-Hybrid
However, there are some perfectly identifiable methods that the community has routinely used, such as the Quantum Approximate Optimization Algorithm (QAOA, [46]), the Variational Quantum Eigensolver (VQE, [47]), or QBSolv [48]. Another interesting and recognizable scheme is Kerberos [49], which is a concretization of the above-described D-Wave Hybrid Framework. More specifically, Kerberos is a reference hybrid workflow comprised of three methods iteratively running in parallel: two classical methods, Simulated Annealing and Tabu Search, and a quantum one that uses the Quantum Processing Unit of D-Wave.
@@
Firstly, NL-Hybrid stands out because it allows for the definition of variables in more advanced formats than those considered in the methods previously included in HSS. More specifically, NL-Hybrid excels with decision variables that embody common logic, such as ordering permutations or subsets of options. For instance, in routing problems like the TSP, a permutation of variables indicates the sequence in which nodes are visited. Similarly, in the KP, the variables representing items to be stored can be categorized into two distinct groups: packed and unpacked.
@@
It is worth mentioning at this point that in the field of optimization, whether by means of classical or quantum systems, the performance of a solver is closely tied to the size of the solution space of the problem at hand. Generally, algorithms perform better in smaller solution spaces. Therefore, employing decision variables that act as implicit constraints is an effective way to reduce the solution space and the complexity of the problem. For instance, using list(number_variables) to represent a TSP implicitly ensures that no nodes are visited more than once along the route. For this reason, the use of the aforementioned decision variables is a significant advantage for NL-Hybrid.
Finally, NL-Hybrid permits linear, quadratic, inequality, and equality constraints, expressed even arithmetically. This aspect represents a significant contribution compared to other well-established hybrid solvers (with the exception of CQM-Hybrid, which also includes this feature).
@@ III Implementation Details
It should be noted that, while the implementation of MCP was done from scratch, those for TSP and KP are slight adaptations of the open-source code published by D-Wave.222https://github.com/dwavesystems/dwave-optimization/blob/main/dwave/optimization/generators.py.
@@ III-A Traveling Salesman Problem
Regarding the objective function, it can be mathematically formulated as follows:
Which is implemented as shown in the following snippet:
@@ III-B Knapsack Problem
Finally, the objective function can be mathematically formulated as follows, being the size of the items set:
Which can be implemented as follows:
@@ III-C Maximum Cut Problem
Furthermore, the objective function can be mathematically formulated as:
Being implemented as represented in the following Python code snippet:
@@ III-D Executing the problem
Once the problem has been modeled following the steps described above, it is ready to be submitted to NL-Hybrid. This process is carried out in the same way as the rest of the HSS portfolio, as can be seen in the following snippet:
@@ IV-A Benchmark Description
- 1.
2.
KP, for which each case is named sX_Y, where X is the number of items and Y is a suffix to distinguish the set of instances with the same X. All instances have been obtained from the KPLib benchmark333https://github.com/likr/kplib, described in [53].
3.
@@ IV-B Experimental Setting
Using the benchmark described above, the main objective is to analyze the performance of the NL-Hybrid, and to compare the results obtained by this method with those obtained by three D-Wave-based counterparts: the full-quantum QPU, BQM-Hybrid and CQM-Hybrid. Our main motivation for choosing these counterparts is to measure the main contribution that NL-Hybrid provides in relation to previously available methods.
Regarding parameterization, the default values have been used for all the solvers for the sake of fairness. For hybrid solver versions, v2.2, v1.12, and v1.1 have been used for BQM-Hybrid, CQM-Hybrid, and NL-Hybrid, respectively. In relation to the QPU, Advantage_system6.4 has been resorted to using the common forward annealing process. Finally, for building the QUBOs introduced as input in both QPU and LeapBQM, Qiskit v0.6.0 open libraries have been employed for the three optimization problems444https://qiskit.org/ecosystem/optimization/apidocs/qiskit_optimization.applications.html.
@@ IV-C Results
In order to obtain representative results, all the outcomes presented in this section have been obtained after 10 independent runs per method and instance. Regarding TSP, we show in Figure 2 the average of the best solution reached by each solver. It should be noted that the metric used to represent the quality of the solutions is the approximation ratio. This metric has been calculated taking as reference the known optimal solution, which have been obtained from [51] and [52].
@@
A simple glance at both figures is enough to detect the superiority of the NL-Hybrid solver over its competitors. To verify this improvement, two statistical tests have been performed with the results shown in Figure 2 and Figure 3.
First, we applied Friedman’s nonparametric test to determine if there are significant differences among NL-Hybrid, CQM-Hybrid, and BQM-Hybrid. It should be noted that QPU has been left out of all these tests as it does not provide outcomes for all datasets. The results of this test are presented in Table III. More specifically, on the left-hand side of the table, we show the results referring to the best results obtained, while on the right-hand side, we show those concerning the complete sample-set mean.
@@
Finally, Figures 7 and 8 present the results for MCP, where the conclusions differ significantly from the previous ones. A glance at the results shows that NL-Hybrid only outperforms the QPU, offering considerably lower performance compared to BQM-Hybrid and CQM-Hybrid, both of which demonstrate outstanding suitability for this problem. To assess if the differences between NL-Hybrid and the other algorithms are statistically significant, we used the Wilcoxon rank-sum test. The results are included in Table VII. Each cell displays the two metrics considered (best solution found and average of the sample set) using one of the following symbols: “” indicates that NL-Hybrid has produced better results than the algorithm in the column with 99% confidence, and “” denotes that the algorithm in the column is statistically superior to NL-Hybrid.
@@
Thus, as can be observed in Table VII, NL-Hybrid significantly outperforms QPU in both metrics, but the superiority of BQM-Hybrid and CQM-Hybrid is also statistically significant for both metrics. However, this performance does not detract an iota of value from NL-Hybrid. Although these results seem counterintuitive compared to those of TSP and KP, they are consistent with what was expressed in Section II-A. That is, depending on the type of problem and the coding of its decision variables, NL-Hybrid may not always be the most suitable algorithm. Specifically, as shown through this experimentation and supported by D-Wave [55], for problems primarily involving binary variables or integers with low ranges, CQM-Hybrid or BQM-Hybrid are more appropriate. This situation opens up a wide range of future work to be carried out around the NL-Solver and its applicability to industrial problems, which we detail in the following section.

### v3

@@ I Introduction
The emergence of quantum technology is expected to have a significant impact on a number of industries. The field of Quantum Computing (QC), which leverages the principles of quantum mechanics to process information, is constantly making advances to connect quantum processing with practical use cases. Thus, QC has advanced significantly in recent years, mainly due to the rapid development of technology and advancements in its democratization, understanding it as the process of making QC accessible to a broader spectrum of individuals and communities [1, 2]. As a result, QC has facilitated the development of various proofs of concept across multiple sectors as finance [3], energy [4], and logistics [5].
Historically, the idea of a quantum computer can be traced back to the works of Benioff [6] and Feynman[7], where they proposed a quantum mechanical implementation of a Turing machine and the idea of simulating a quantum system with another quantum system, respectively. Currently, there are two types of real quantum devices that we can differentiate between: gate-based quantum computers and quantum annealers. A gate-based system, on the one hand, employs basic quantum circuit operations on qubits, which are akin to the classical operations on regular bits and may be combined in any order to create algorithms. This version is often referred to as a universal quantum computer. On the other hand, a quantum annealer is based on adiabatic computation, wherein an initial, easily prepared Hamiltonian is gradually and continuously evolved from its ground state to the ground state of a final, problem-specific Hamiltonian. If the evolution is slow enough, the adiabatic theorem guarantees that the system remains in the ground state during the whole computation. In quantum annealers, the adiabatic theorem is intentionally relaxed, allowing the system to evolve more rapidly than the adiabatic limit would dictate. As a result, transitions to higher energy states usually occur during the evolution; however, alternative methods to achieve adiabaticity have been proposed [8, 9]. Although this model of computation is also universal [10], D-Wave quantum annealer is based on an Ising Hamiltonian, restricting the types of problems that can be run on the system. This type of quantum annealer is, however, well suited to solve combinatorial optimization problems [11].
@@
Despite all the progress made in the field, quantum computers are still in their infancy in contrast to classical computers, which have been developed for decades and are therefore highly advanced. Thus, quantum devices are currently unable to efficiently solve real-world problems mainly because of the small number of qubits and their unstable nature. Phenomena such as decoherence time, noise and information loss, in absence of error correction protocols, impact in the performance of the computation. Besides, there are other roadblocks such as quantum gate fidelity and gate noise. At this time, the number of qubits for both universal quantum computers and quantum annealers is around three orders of magnitude (thousands of qubits) [12, 13]. However, this number needs to be much greater in order for the technology to become truly useful for real industrial use cases. In any case, there are factors other than the qubit count that influence the practical capabilities of quantum computers, such as the connectivity topology or the previously mentioned operational fidelity and decoherence time.
@@
Besides, hybrid algorithms have its own limitations. For example, since quantum and classical hardware are often physically separated and they need to share information, latency is introduced in the workflow [17]. Another roadblock is the possibility of barren plateaus in the optimization landscape, making it difficult for classical optimizers to converge [18], [19]. Also, when the problem at hand exceeds the computational capacity of the quantum processing unit (QPU), determining which subproblems to allocate to the quantum component of the algorithm is a non-trivial task [20], [21]. The choice of methodology is not always straightforward and often depends on the specific characteristics of the problem and the quantum hardware’s limitations, for example its topology.
It is also crucial to note that the design and implementation of quantum algorithms could necessitate a high degree of subject matter expertise. This intricacy could be a hindrance for researchers who lack a sufficient background in disciplines like physics or quantum mechanics. In order to remove that barrier and make QC more accessible, various frameworks and programming languages are proposed, such as Silq [22], Eclipse Qrisp [23], or Qiskit [24]. The availability of such frameworks and languages helps to foster the building of a multidisciplinary community focused on QC and helps the field to progress toward new horizons [25].
In this article, we focus on the hybrid algorithms and frameworks proposed by the Canadian company D-Wave Systems. Anyway, it is worth noting that significant efforts have been made from various fronts in the design and implementation of both hybrid techniques and platforms. Probably the most well-known and widely studied hybrid resolution schemes are the Variational Quantum Algorithms (VQA, [26])), oriented towards gate-based quantum computing. The most representative examples of VQA are the Quantum Approximate Optimization Algorithm (QAOA, [27]) and the Variational Quantum Eigensolver (VQE, [28]). In this context, it is also interesting to highlight commercial approaches such as the so-called HybridSolver developed by the German company Quantagonia111https://www.quantagonia.com/hybridsolver. Regarding frameworks and platforms specifically created to facilitate the design, implementation, and execution of hybrid algorithms, notable examples include NVIDIA’s CUDA-Q222https://developer.nvidia.com/cuda-q, IBM Qiskit Runtime333https://docs.quantum.ibm.com/api/qiskit-ibm-runtime/runtime_service and Google’s Cirq444https://quantumai.google/cirq.
Focusing our attention on D-Wave, and as part of the strategy to bring QC to a wider audience, D-Wave-Hybrid-Framework was published in 2018, which, in the words of its creators, is “a general, minimal Python framework to build hybrid asynchronous decomposition samplers for QUBO problems”555https://github.com/dwavesystems/dwave-hybrid. This framework is appropriate for “developing hybrid approaches to combining quantum and classical compute resources”.
@@
As we have summarized systematically in Table I, much research has been carried out in recent years around BQM-Hybrid, DQM-Hybrid, and CQM-Hybrid. Among these methods, BQM-Hybrid is the one most frequently used by the scientific community. This is because it was the first to be implemented and its usage is similar to the D-Wave QPU. It is important to highlight that, precisely for this reason, the BQM-Hybrid is not well-suited for solving real-world problems with many constraints. Consequently, there are few studies where this method has been used to solve problems with realistic aspects [31, 32, 33, 34]. With all this, the BQM-Hybrid has been primarily used as a benchmarking algorithm [35, 36, 37, 38, 39, 40, 41, 42]. That is, as a method employed within an experimentation to measure the performance of other methods. Moreover, BQM-Hybrid has also been used to solve academic problems, such as Tail Assignment Problem [43], Vehicle Routing Problem [44] and the Set Packing Problem [45].
Focusing on the CQM-Hybrid, we see how this advanced solver has a clear orientation towards solving real-world problems. The CQM-Hybrid has been used in a significant number of works to address complex problems with a high number of constraints, and it has been employed in different contexts: to solve a complete problem [46, 47, 48, 35, 39, 49, 50]; as part of a complex resolution pipeline where it is called a single time [51] or iteratively [52]; or to implement mechanisms to improve the performance of existing classical algorithms [53]. Furthermore, the CQM-Hybrid has been scarcely used as a benchmarking algorithm [42, 54].
Regarding the DQM-Hybrid, it is the HSS method that has received the least attention from the community. This is mainly because it is less flexible and powerful compared to CQM-Hybrid, leading researchers to clearly favor the latter. Thus, the DQM-Hybrid has been used in only a handful of papers, either to act as a benchmarking method [39] or to solve clustering and community detection problems [55, 56, 57].
Despite this abundant scientific activity, no work has yet been published on NL-Hybrid. Motivated by the lack of existing research, this paper focuses on describing the newly introduced NL-Hybrid solver. Furthermore, we will conduct an experiment to analyze the performance of this new solver in comparison with BQM-Hybrid, CQM-Hybrid, and D-Wave’s QPU Advantage_system6.4. For these tests, we used a benchmark composed of 45 instances, equally distributed across three combinatorial optimization problems: the Traveling Salesman Problem (TSP, [59]), the Knapsack Problem (KP, [60]), and the Maximum Cut Problem (MCP, [61]). We have chosen these problems because:
They are appropriate for formulation and use in the solvers considered in this study. Furthermore, TSP and KP are well suited to be solved with NL-Hybrid newly introduced variables, while MCP can be easily formulated with traditional binary variables.
Finally, to facilitate the use of this still unexplored solver, we will share details of the implementation to solve these three optimization problems.
This research complements previous work on hybrid algorithms carried out by the authors of this study. In works such as [67, 68], the authors presented ad hoc implemented hybrid algorithms to solve the TSP and its asymmetric variant. Furthermore, in [53, 52, 47, 46], the authors used the CQM-Hybrid to tackle real-world problems in fields such as logistics and industry. Finally, the authors have also conducted research focused on benchmarking different hybrid algorithms, with representative works such as [69, 70, 71]. This paper stands out from all these papers, being the first to work with and experiment on the newly unveiled NL-Hybrid.
@@ II-A Overview of NL-Hybrid
However, there are some well-established methods that the community routinely uses, such as the aforementioned QAOA and VQE, or QBSolv [73]. Another interesting and recognizable scheme is Kerberos [74], which is a concretization of the above-described D-Wave Hybrid Framework. More specifically, Kerberos is a reference hybrid workflow comprised of three methods iteratively running in parallel: two classical methods, Simulated Annealing and Tabu Search, and a quantum one that uses the QPU of D-Wave.
@@
Firstly, NL-Hybrid stands out because it allows for the definition of variables in other additional formats than those considered in the methods previously included in HSS. More specifically, NL-Hybrid excels with decision variables that embody common logic, such as ordering permutations or subsets of options. For instance, in routing problems like the TSP, a permutation of variables indicates the sequence in which nodes are visited. Similarly, in the KP, the variables representing items to be stored can be categorized into two distinct groups: packed and unpacked.
@@
It is worth mentioning at this point that in the field of optimization, whether by means of classical or quantum systems, the performance of a solver is closely tied to its capacity to explore and exploit the whole solution space. That is, the larger the search space, the higher the probability of reaching better solutions, understanding search space as the region of the solution space that the algorithm can access. However, there is the downside that larger spaces are usually computationally expensive to explore and exploit. Therefore, employing decision variables that act as implicit constraints is an effective way to reduce both search and solution spaces and thus, the running time to find a solution. For instance, using list(number_variables) to represent a canonical TSP implicitly ensures that no nodes are visited more than once along the route and also that each node is visited. This results in the absence of infeasible solutions within the solution space, thereby ensuring that this space comprises solely feasible solutions to the problem. This characteristic is not observed with binary encoding. For this reason, the use of the aforementioned decision variables is a significant advantage for NL-Hybrid.
Finally, NL-Hybrid natively permits nonlinear (linear, quadratic, and higher order) inequality and equality constraints, expressed even arithmetically. This aspect represents a significant contribution compared to other well-established hybrid solvers. For comparison, BQM-Hybrid accepts linear soft constrains as penalty models, while CQM-Hybrid works with linear and quadratic constrains natively, although it is also possible to implement them as soft constraints.
@@ II-B Structure and Workflow of NL-Hybrid
Lastly, it should be noted that the NL-Hybrid solver is proprietary. Consequently, further technical details are not available to the general public. For additional information on this method, we refer interested readers to the D-Wave report on the HSS portfolio [29].
@@ III Implementation Details
It should be noted that, while the implementation of MCP was done from scratch, those for TSP and KP are slight adaptations of the open-source code published by D-Wave666https://github.com/dwavesystems/dwave-optimization/blob/main/dwave/optimization/generators.py.
@@ III-A Traveling Salesman Problem
Regarding the objective function, which must be minimized, it can be mathematically formulated as follows, being a list representing a feasible TSP route:
Which is represented in code as shown in the following snippet:
@@ III-B Knapsack Problem
Finally, the objective function, which must be maximized natively, can be mathematically formulated as follows, being a feasible set of items, and the size of this set:
Which can be represented in code as follows:
@@
It is worth noting that NL-Hybrid method only allows for the minimization of the objective function. This is why problems natively designed to maximize an objective, such as the KP, are handled by adding a negative sign to the function, as seen in the code above. This situation also occurs with the MCP.
@@ III-C Maximum Cut Problem
Furthermore, the objective function, which must be maximized natively, can be mathematically formulated as follows, being a feasible binary solution to the MCP:
Which can be made explicit in code using the following this Python snippet:
@@ III-D Executing the problem
Once the problem has been modeled following the steps described in the previous subsections, it is ready to be submitted to NL-Hybrid. This process is carried out in the same way as the rest of the HSS portfolio, as can be seen in the following snippet:
@@
The code above produce an outcome as we represent in the following code snippet, centered on a TSP example composed of 9 nodes. {python} solution - objective function [7 8 6 4 2 3 0 1 5] - 2134.0 [1 5 7 8 6 4 2 3 0] - 2134.0 [1 0 3 2 4 6 8 7 5] - 2134.0 [4 6 8 7 5 1 0 3 2] - 2134.0 [3 2 4 6 8 7 5 1 0] - 2134.0 …
@@ IV-A Benchmark Description
- A.
B.
KP, for which each case is named sX_Y, where X is the number of items and Y is a suffix to distinguish the set of instances with the same X. All instances have been obtained from the KPLib benchmark777https://github.com/likr/kplib, described in [77].
C.
@@ IV-B Experimental Setting
Using the benchmark described above, the main objective is to analyze the performance of the NL-Hybrid, and to compare the results obtained by this method with those obtained by three D-Wave-based counterparts: the QPU, BQM-Hybrid and CQM-Hybrid. Our main motivation for choosing these counterparts is to measure the main contribution that NL-Hybrid provides in relation to previously available methods. More specifically, the BQM-Hybrid has been employed in these tests because it is the most widely employed HSS solver and its input format is similar to that of the QPU. Additionally, the CQM-Hybrid has been used because it is the most advanced algorithm in the portfolio and is easy to use when modeling the problems addressed. Finally, the DQM-Hybrid has been set aside due to the little attention it has received in the literature and the complexity of its use compared to the other solvers within HSS.
Regarding the QPU, the Advantage_system6.4 device has been used, which is the most recent from D-Wave at the time this work was written. This computer features 5,616 qubits and more than 35,000 couplers arranged in a Pegasus topology. Similar to hybrid solvers, the QPU has been accessed through the D-Wave Leap cloud service, and the common forward annealing process has been executed, which consists of the following main steps [79]: i) convert QUBO into a graph, ii) minor graph embedding, iii) QPU initialization, iv) annealing, v) readout, and vi) postprocessing. In Figure 2, we graphically represent this process to facilitate its understanding.
In relation to the parameterization, the default values have been used for all the solvers for the sake of fairness. For hybrid solver versions, v2.2, v1.12, and v1.1 have been used for BQM-Hybrid, CQM-Hybrid, and NL-Hybrid, respectively. For the QPU, the default parameters have also been used.
Finally, Qiskit_Optimization888https://qiskit.org/ecosystem/optimization/apidocs/qiskit_optimization.applications.html v0.6.1 open libraries have been employed to aid in solving the three optimization problems using the QPU and BQM-Hybrid. These libraries have been utilized because they enable tasks such as data reading, solution decoding, result evaluation, and, most importantly, the automatic construction of the QUBOs. More specifically, to build the QUBO formulations, these classes relax the constraints using a penalty model, with the coefficient being automatically estimated. Finally, it is worth noting that Qiskit_Optimization does not present any incompatibility with any versions of Qiskit, such as the current v1.2.
@@ IV-C Results
In order to obtain representative results, all the outcomes presented in this section have been obtained after 10 independent runs per method and instance. Regarding TSP, we show in Figure 3 the average of the best solution reached by each solver. It should be noted that the metric used to represent the quality of the solutions is the approximation ratio. This metric has been applied to all the methods executed, and it is particularly suitable for measuring the distance between the obtained outcomes and known optimal solutions in terms of the objective function cost. Thus, the approximation ratio is calculated by using the cost of an obtained solution and the optimal value of the problem being solved. The optimal values for each instance have been obtained from [71] and [76].
@@
A simple glance at Figure 3 and Figure 4 is enough to detect the superiority of the NL-Hybrid solver over its competitors. To verify this improvement, two statistical tests have been performed with the results shown in these figures.
First, we applied Friedman’s non-parametric test to determine if there are significant differences among NL-Hybrid, CQM-Hybrid, and BQM-Hybrid. It should be noted that QPU has been left out of all these tests as it does not provide outcomes for all datasets. The results of this test are presented in Table III. Specifically, the average ranking value returned by the Friedman non-parametric test is presented for each of the compared algorithms. A lower rank indicates better performance. Furthermore, on the left-hand side of Table III, we show the results referring to the best results obtained, while on the right-hand side, we show those concerning the complete sample-set mean.
@@
Finally, Figures 8 and 9 present the results for MCP, where the conclusions differ significantly from the previous ones. A glance at the results shows that NL-Hybrid only outperforms the QPU, offering considerably lower performance compared to BQM-Hybrid and CQM-Hybrid, both of which demonstrate outstanding suitability for this problem. To assess if the differences between NL-Hybrid and the other algorithms are statistically significant, we used the Wilcoxon rank-sum test. The results are included in Table VII. Each cell displays the two metrics considered (best solution found and average of the sample set) using one of the following symbols: “” indicates that NL-Hybrid has produced better results than the algorithm in the column with 99% confidence, and “” denotes that the algorithm in the column is statistically superior to NL-Hybrid.
@@
Thus, as can be observed in Table VII, NL-Hybrid significantly outperforms QPU in both metrics, but the superiority of BQM-Hybrid and CQM-Hybrid is also statistically significant for both metrics. However, this performance does not detract from the value of NL-Hybrid. Although these results seem counterintuitive compared to those of TSP and KP, they are consistent with what was expressed in Section II-A. That is, depending on the type of problem and the coding of its decision variables, NL-Hybrid may not always be the most suitable algorithm. Specifically, as shown through this experimentation and supported by D-Wave [80], for problems primarily involving binary variables or integers with low ranges, CQM-Hybrid or BQM-Hybrid are more appropriate. This situation opens up a wide range of future work to be carried out around the NL-Solver and its applicability to industrial problems, which we detail in the following section.